OpenStudy (anonymous):

A ball is thrown downwards with speed 5m/s from a top a 20m high building. a. How long does it take for the ball to hit the ground? b. How fast is it moving when it hits the ground?

OpenStudy (jt950):

(a) S=u t+(1/2) gt^2 20=5 t+(1/2) (9.8) t^2 t=1.57 second (b) V=u+gt =5+(9.8)(1.57) =20.42m/s

OpenStudy (anonymous):

thnks

OpenStudy (anonymous):

what are the variables S , u

OpenStudy (jt950):

S is distance travelled u is initial velocity v is final velocity

OpenStudy (jt950):

there is 1 more equation for this v^2=u^2+ 2as

OpenStudy (anonymous):

so which one should i use

OpenStudy (jt950):

All you can use depend on the situation. In this situation , S,u,g is given, you decide what to use :)

OpenStudy (anonymous):

ok thnkds a lot

OpenStudy (anonymous):

im still confused as to calculating the first part the units are messing me up and also you have t and t^2

OpenStudy (jt950):

use SI unit m, m/s , s , m/s^2 t^2 and t use quadratic equation to solve you get 1 negative(rejected) 1 positive(1.57)